A waterfall is 30 meters high, and 10000 kg of water flows over it per second. How much power does this flow generate? If all this power is converted to electricity, how many 100-watt bulb can it supply?


1
Expert's answer
2021-01-23T14:08:10-0500

a) By the definition of the power, we have:


"W=\dfrac{W}{t}=\dfrac{mgh}{t},""W=\dfrac{10000\ kg\cdot 9.8\ \dfrac{m}{s^2}\cdot30\ m}{1\ s}=2.94\cdot10^6\ W."

b) Finally, we can find the number of 100-watt bulbs that can supply at this power:


"N=\dfrac{2.94\cdot10^6\ W}{100\ W}=29400\ bulbs."